OpenStudy (anonymous):

Refer to the table:Write and Solve an equation to find each quantity. 1tbs= 3tsp 1c=16tbs 1c=8fl. oz. 1 pint=2c. 1qt=2pints 1gal.=4qts. Write an equation to show how to solve the number of cups in 96 teaspoons.

OpenStudy (stacey):

\[\frac{ 96 tsp }{1 }*\frac{ 1 tbs }{3 tsp }*\frac{ 1 c }{ 16 tbs }\]

OpenStudy (anonymous):

Can you explain that?

Directrix (directrix):

@Stacey OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution.

OpenStudy (stacey):

If two values are equal, then their ration is equivalent to 1. Multiplying by 1 does not change the value of a fraction. So we start with the 96 tsp that you are given. Whole numbers can be put over 1 to make them into a fraction.  We want to cancel out the teaspoons in the numerator, so we select an equality that has teaspoons (1 tbs = 3 tsp) and make a fraction out of it with the teaspoons in the denominator. Since we do not want tbs, we find an equality with tbs so we can cancel them. We have two choices. Since we want to end up with cups, we choose 1 c = 16 tbs.